### v3

@@ 4.3.2 Performance result SATD categorization
NLP Annotation tags Annotation tags Annotation tags
ML SMO NB SMO
@@
Transformers BERT BERT BERT
The best-performing algorithms for each approach based on the different SATD types is presented in Table 10. In both the SATD identification and categorization tasks, BERT consistently emerges as the best-performing model. For SATD identification, BERT’s high median F1-score, tested across a large number of experiments (n=69), showcases its strong reliability and generalizability. Its low variance and minimal outliers further demonstrate its consistent and stable performance across different datasets. Leveraging its advanced Transformer-based architecture, BERT excels at capturing the context and nuances in SATD instances, making it highly effective in balancing precision and recall.
Similarly, in the SATD categorization task, BERT outperformed other models, achieving the high median F1-score with a significant number of experiments (n=47). Its minimal variance and absence of outliers reinforce its reliability and consistency in this task as well. Altogether, these factors position BERT as the most robust and effective model for both SATD identification and categorization, making it the top choice in each task.
Answer to RQ3 BERT (a Transformers-based model) emerges as the top performer in both SATD and non-SATD identification and the categorization of specific SATD types, providing balanced precision, recall, and F1 scores. SMO (an ML model) and GGNN (a DL model) also show strong performance within their respective domains, with SMO excelling in precision and F1 score, while GGNN leads in identifying specific SATD types. Overall, Transformer models, particularly BERT, represent the most advanced and effective approach to SATD detection.
@@ 4.4 Results - RQ4: Proposed tools for SATD detection
To address this research question, we conducted a thorough review of existing tools developed over the past decade. While the detection and management of SATD have been extensively studied, there remains a noticeable gap in translating research into practical tools for developers. This section examines tools specifically designed for SATD detection, offering insights into advancements in automation and machine learning, as well as the diverse approaches used to address various forms of technical debt.
@@
Table 11 presents a collection of tools proposed for SATD identification across different software projects. This compilation reflects the evolution of tool development over time, highlighting an increasing emphasis on automation, machine learning integration, and varied strategies for managing different types of technical debt. Below, we describe these tools and their significance in the broader context of SATD identification, emphasizing key trends that have emerged over the last decade.
Table 11: SATD detection tools proposed
@@
Name Reference Year Task Type Description Status
eXcomment No. 36, 48 2015 Identification A tool designed to parse Java source code and fetch code comments to identify SATD or non-SATD Broken link
SATD Detector No. 55 2018 Identification A Java library and an Eclipse plug-in. The tool automatically identifies whether comments are SATD or non-SATD and integrates with an IDE to help manage these comments Obsolete
DebtHunter No. 64 2021 Identification, Categorization A machine learning-based approach for detecting self-admitted technical debt Available
FixMe No. 72 2021 Categorization A GitHub bot that is developed to detect, monitor, and notify developers about On-hold SATD in their repositories Broken link
SATDBailiff No. 67 2022 Identification A tool designed to mine, identify, and track SATD Available
A browser extension No. 74 2022 Categorization A browser extension utilizes a developed ML model to automatically categorize the SATD types (i.e., documentation, code, defect, test, design, build, architecture, versioning, usability, and requirement debt) of a rOpenSci R package Not available
@@
DebtViz No. 90 2023 Categorization A tool for SATD detection, classification, visualization, and monitoring in a single platform to categorize code/design, documentation, test, and requirement debt types Available
The tools proposed for SATD identification reflect a diverse range of approaches to addressing technical debt in software projects. eXcomment 36, 48, and SATD Detector 55 focus on identifying both SATD and non-SATD comments, with eXcomment integrating into an IDE for real-time detection. A browser extension 74 leverages machine learning to detect various SATD types, including documentation and usability debt, highlighting the potential of ML for broader SATD detection. SATDBailiff 67 provides comprehensive tracking and monitoring of SATD over time, while FixMe 72 is a GitHub bot specifically designed to alert developers about unresolved, “on-hold SATD" in repositories. Lastly, DebtViz 90 offers a more complete solution, combining SATD detection, classification, visualization, and monitoring in one platform, emphasizing the need for visual tools that make technical debt more understandable to stakeholders. Together, these tools demonstrate an increasing focus on automation, monitoring, and comprehensive SATD management, though the overall number of tools remains limited relative to the research interest in the field.
@@
This trend of limited tool development is further highlighted when considering that, despite the SLR encompassing 68 relevant studies, only 7 articles proposed specific tools for SATD identification. This disparity points to a concerning gap between the significant research interest in SATD and the actual development of practical, automated solutions that can assist developers in managing technical debt. While a range of tools has been introduced, the overall lack of available tools indicates that much of the research has yet to translate into actionable, industry-ready solutions. Addressing this gap will be critical for advancing the practical management of technical debt in software projects, moving from theoretical frameworks toward robust, scalable automation.
@@
Answer to RQ4 We identify several tools proposed for SATD detection, with a focus on automation, machine learning integration, and addressing different types of technical debt. However, despite significant research interest, only a few tools have been developed, revealing a gap between theoretical advancements and the availability of practical, automated solutions for managing technical debt in real-world software projects.
@@ 4.5 Results - RQ5: Primary challenges associated with SATD detection
We identified nine key challenges related to SATD identification extracted from the relevant studies, as summarized in Table 12, with relevant studies emphasizing each challenge. These challenges were determined by examining sections commonly located in the ‘Discussion’ or ‘Threat to Validity’ sections of the papers reviewed. We extracted all discernible challenges, whether implicitly or explicitly mentioned, associated with SATD detection. Each challenge is explained in this section and further elaborated on below.
Table 12: Papers highlighting key challenges. Studies that fall into more than one challenge are highlighted with square brackets
@@
1 Error-proneness dues to manual analysis No. 8, 38,[37, 36, 50, 53, 59, 39, 49]
2 Imbalanced data No. [54, 77, 78, 40, 79, 60, 61, 62, 64, 81, 82, 83, 45, 85, 86, 88, 68, 93, 94, 96, 71], 56, 63, 80, 66, 89, 46, 47, 92, 75, 95, 97
3 Generalizability No. [50, 53, 59, 39, 37, 55, 77, 40, 60, 61, 64, 78, 81, 45, 85, 86, 88, 68, 93, 94, 96, 49, 99, 71], 58, 74, 101, 87, 65, 69
@@
Manual Analysis
The first challenge is the dependence on manual analysis, which introduces the potential for human error and inconsistency 38. Since the process relies on human judgment, it becomes prone to inconsistencies in labeling, as different individuals may interpret and categorize SATD differently 36. This subjectivity in analysis can lead to varying outcomes in tasks like pattern recognition and feature engineering, making the process less reliable.
The variability in human interpretation introduces a lack of uniformity in the labeled datasets, which can severely impact the effectiveness of machine learning models that rely on this data for training. When models are trained on datasets that contain inaccuracies or inconsistencies, their predictive accuracy and generalization capabilities suffer. This challenge becomes even more pronounced when scaling up to larger datasets, where consistent and precise labeling is critical to maintaining the quality of the training data. Without robust and reliable analysis, the models may struggle to handle the increased complexity of larger data, leading to reduced performance and misidentification of SATD instances.
@@
In most software projects, SATD comments represent only a small portion of the total code comments, resulting in a significant class imbalance. This imbalance can adversely affect the performance of SATD identification models, as the models may tend to focus on non-SATD comments, which dominate the dataset. As a result, the models may struggle to accurately identify the relatively rare SATD instances, often prioritizing the majority class and overlooking the minority class 66.
This imbalance frequently leads to a high number of false negatives 105, where the model fails to detect SATD instances, undermining its effectiveness. Since machine learning models generally find it difficult to learn patterns from underrepresented classes 106, the scarcity of SATD examples makes it challenging for these models to develop a robust understanding of SATD characteristics. Consequently, the model’s ability to identify and classify SATD instances diminishes, particularly when faced with a large volume of non-SATD items. Addressing this imbalance is critical to improving the precision and reliability of SATD identification models, as it ensures that the models are not biased toward detecting the majority class while neglecting the crucial SATD instances.
@@
One of the primary challenges associated with the identification of SATD is the issue of generalizability. Generalizability refers to the ability of SATD identification models to maintain their accuracy and effectiveness when applied across different software projects and domains. This challenge arises due to the diverse nature of software projects, which often exhibit significant variations in coding practices, domain-specific terminologies, and architectural designs 44.
Models trained on data from one specific project tend to be tailored to the characteristics of that project, limiting their performance when applied to other projects with different practices. For example, the way SATD is expressed in comments may differ across projects due to differences in the development team’s language, domain jargon, or code style conventions 58, 39. These project-specific factors can result in models that are highly effective within their training environment but struggle to generalize to other contexts, especially when faced with unfamiliar coding practices or new domains.
@@
Another primary challenge in the SATD identification process, particularly in machine learning and deep learning-based models, is the lack of explainability. This issue arises because many advanced models, such as neural networks and deep learning architectures, function as “black boxes," meaning their internal decision-making processes are not easily interpretable by users 107. These models generate predictions without providing transparent explanations about the underlying reasoning for classifying an instance as SATD, non-SATD, or even specific types of SATD.
@@
False negatives are another concern in SATD detection. They occur when SATD items are incorrectly classified as non-SATD, leading to undetected technical debt in the codebase. This happens when models fail to identify instances where developers have either implicitly or explicitly admitted technical debt 38. As a result, undetected technical debt can linger in the system, leading to long-term code degradation, higher maintenance costs, and an increased risk of introducing more bugs and system failures 12, 108.
@@
Another primary challenge in the context of machine learning and supervised models-based models is the issue of reproducibility 44, 67. These models often involve complex architectures with numerous parameters that must be fine-tuned for optimal performance. However, many studies fail to provide sufficient details on the hyperparameters, training configurations, or the specific processes used to achieve their results. Additionally, the lack of publicly available code exacerbates the reproducibility problem, as other researchers are unable to replicate the exact experimental setup. This makes it difficult to reproduce results across different studies and datasets, raising concerns about the reliability and validity of these approaches.
@@
Another challenge associated with SATD detection is the high computational costs of many advanced models, particularly deep learning approaches. These models require significant resources during training, including high-performance GPUs, large memory capacities, and extended computation time. This high computational cost creates several obstacles for researchers and practitioners who aim to implement or improve SATD identification methods.
Additionally, the extended computation time required for training deep learning models presents a challenge in terms of practicality. SATD identification models often need to be retrained or fine-tuned to adapt to new projects or datasets, and the time-consuming nature of this process can hinder agile software development practices. In scenarios where rapid feedback is essential, such as continuous integration and deployment environments, the lengthy training times of deep learning models can become a bottleneck, delaying the detection of technical debt and the subsequent mitigation efforts.
@@
Furthermore, the overlap of features among different types of technical debt, such as code debt, design debt, and requirement debt, presents another challenge 66. This overlap makes distinguishing between different debt types difficult, leading to potential misclassification. The ambiguity in distinguishing between SATD and other debt types can reduce the accuracy of identification models. If features that are common to multiple debt types are not carefully handled, models may struggle to correctly identify SATD.
@@
Finally, tagging and comment relevance is a critical challenge in identifying SATD. Developers often use tags such as “TODO, “FIXME," or “XXX" in code comments or other textual artifacts to mark areas that require attention or indicate temporary fixes. While these task-related tags are commonly employed as indicators of SATD, they are not always relevant to technical debt. In many cases, tags like “TODO" or “FIXME" can represent a simple reminder or a non-technical debt issue, thereby creating noise in the identification process. Moreover, not all SATD instances are explicitly marked by such tags, further complicating the identification process.
Answer to RQ5 We identified nine major challenges in detecting SATD, including issues like manual analysis errors, imbalanced data, lack of generalizability, and explainability in advanced models. Other challenges include false negatives, reproducibility risks, computational costs, overlapping features among debt types, and the relevance of tagging and comments, all of which hinder effective SATD identification.

### v4

@@ 4.3.2 Performance result SATD categorization
Heuristic Annotation tags Annotation tags Annotation tags
ML Auto-sklearn Auto-sklearn Auto-sklearn
@@
Transformer BERT BERT BERT
The best-performing algorithms for each technique and SATD task are summarized in Table 10. Overall, Transformer-based models consistently deliver the strongest reported results across both SATD identification and categorization. For SATD identification, DistilRoBERTa yields the highest precision, recall, and F1 score among Transformer variants in our synthesis, indicating strong and balanced effectiveness in binary classification settings. ML models, particularly Random Forest, also show competitive performance for identification, especially in precision and F1, while among DL models BiLSTM achieves the strongest overall balance between precision and recall. For SATD categorization, BERT remains the best-performing Transformer model, achieving the highest precision, recall, and F1 score, while Auto-sklearn and GGNN lead the ML and DL groups, respectively. These outcomes reflect trends reported across the literature rather than a definitive ranking, as performance is strongly influenced by dataset quality, annotation practices, and evaluation protocols.
Answer to RQ3 Overall, Transformer-based models achieve the strongest reported effectiveness across SATD detection tasks. In SATD identification, DistilRoBERTa attains the highest precision, recall, and F1 score in the synthesized results, while in SATD categorization BERT is the best-performing model across all three metrics. Although several ML (e.g., RF and Auto-sklearn) and DL models (e.g., BiLSTM and GGNN) also achieve strong results in specific settings, their performance is generally less consistent than Transformer-based techniques across the heterogeneous datasets and evaluation protocols used in the primary studies.
@@ 4.4 Results - RQ4: Proposed tools for SATD detection
To address this research question, we conducted a comprehensive review of existing tools developed over the past decade. While SATD detection and management have been extensively studied, there remains a significant gap in translating research findings into practical, developer-friendly tools. This section focuses on tools specifically designed for SATD detection, providing insights into advancements in automation and the various techniques used to address different forms of technical debt.
In this SLR, we classify an approach as a tool only when the authors explicitly propose a prototype or a ready-to-use system, rather than merely providing source code. To be considered a valid contribution to SATD detection, we regard a tool as a system designed for external use, meaning it can be executed with minimal setup and without requiring significant modifications and must be deployable and practically usable 115, 116. While some studies provide source code to support their experiments, the availability of source code alone is not sufficient to qualify an approach as a tool. Additionally, we include studies in which a tool is described and a replication package is provided, even if the authors do not explicitly refer to it as a tool or if the provided link is no longer valid, as long as the paper provides sufficient evidence that the tool was originally intended for SATD detection.
@@
Table 11 presents a collection of SATD detection tools proposed in different software projects. This compilation reflects the evolution of tool development over time, highlighting an increasing emphasis on automation, ML integration, and varied strategies for managing different types of technical debt. The following discussion examines the significance of these tools within the broader context of SATD detection and identifies key trends that have shaped their development over the past decade.
Table 11: SATD detection tools proposed.
@@
Name Year Technique Automation Level Artifact Task Type Description Status
BEACon-TD 109 2025 Transformer Automated Issue Identification, Categorization DistilRoBERTa-based tool that classifies software issues as technical debt and further categorizes them into 13 types using a two-stage ensemble of binary classifiers Available
@@
DebtViz 98 2023 DL Automated Code comment & Issue Categorization A tool for SATD detection, classification, visualization, and monitoring in a single platform to categorize code/design, documentation, test, and requirement debt types Available
SATDBailiff 74 2022 ML Automated Code comment Identification A tool designed to mine, identify, and track SATD Available
A browser extension 83 2022 ML Automated Issue Categorization A browser extension utilizes a developed ML model to automatically categorize the SATD types (i.e., documentation, code, defect, test, design, build, architecture, versioning, usability, and requirement debt) of an rOpenSci R package Not available
DebtHunter 72 2021 ML Automated Code comment Identification, Categorization A machine learning-based technique for detecting self-admitted technical debt Available
FixMe 79 2021 ML Automated Code comment Categorization A GitHub bot that is developed to detect, monitor, and notify developers about On-hold SATD in their repositories Broken link
SATD Detector 64 2018 ML Automated Code comment Identification A Java library and an Eclipse plug-in. The tool automatically identifies whether comments are SATD or non-SATD and integrates with an IDE to help manage these comments Obsolete
eXcomment 47, 58 2015 Heuristic Semi-Automated Code comment Identification A tool designed to parse Java source code and fetch code comments to identify SATD or non-SATD Broken link
The proposed tools for SATD identification employ diverse techniques and varying levels of automation to address technical debt in software projects. eXcomment 47, 58 utilizes a heuristic approach with semi-automated functionality to parse Java source code comments and detect SATD. In contrast, SATD Detector 64 applies ML in a fully automated manner to identify and classify SATD derived from code comments within an IDE, although it has since become obsolete.
Advancing the role of ML, tools such as DebtHunter 72 and FixMe 79 implement automated SATD detection and categorization. DebtHunter employs ML to detect SATD, while FixMe functions as a GitHub bot that actively monitors and notifies developers about unresolved On-hold SATD.
SATDBailiff 74 further extends SATD management by introducing an automated tool designed to mine, track, and monitor SATD occurrences over time. Another approach is demonstrated by a browser extension 83, which leverages an ML model to automatically categorize different SATD types, such as documentation, usability, and requirement debt. Lastly, DebtViz 98 adopts DL to provide an automated solution that integrates detection, classification, visualization, and monitoring of various SATD types within a single platform. More recently, BEACon-TD 109 introduces a Transformer-based solution that detects SATD and classifies it into 13 specific types using a two-stage ensemble of binary classifiers.
@@
This trend of limited tool development is further highlighted when considering that, despite the SLR encompassing 74 relevant studies, only 8 articles proposed specific tools for SATD identification. This disparity highlights a concerning gap between the substantial research interest in SATD and the development of practical, automated solutions to assist developers in managing technical debt. While a range of tools has been introduced, their limited availability indicates that much of the research has yet to translate into actionable, industry-ready solutions. Addressing this gap will be critical for advancing the practical management of technical debt in software projects, moving from theoretical frameworks toward robust, scalable automation.
@@
Answer to RQ4 We identify several tools proposed for SATD detection, with a focus on automation, ML integration, and addressing different types of technical debt. However, despite significant research interest, only a few tools have been developed, revealing a gap between theoretical advancements and the availability of practical, automated solutions for managing technical debt in real-world software projects.
@@ 4.5 Results - RQ5: Primary challenges associated with SATD detection
We identified nine key challenges related to SATD identification from the relevant studies, as summarized in Table 12, and each study emphasizes one of these challenges. The extraction process focused on statements reported in the ‘Discussion’ or ‘Threats to Validity’ sections of the reviewed papers. All discernible challenges, whether explicitly or implicitly mentioned, were collected for further analysis.
The grouping of challenges was also performed using a bottom-up thematic analysis. Each extracted challenge statement was coded and then clustered into higher-level categories based on thematic similarity, following common practices in prior SE literature reviews. The first author conducted the initial coding and grouping, while the second author independently cross-checked a 15% random sample of the classifications using the shared spreadsheet. Any disagreements were resolved through collaborative discussion until a consensus was reached. Only one case required clarification, indicating that disagreements were minimal and consistently resolvable. Each resulting challenge category is explained in this section and further elaborated on below.
Table 12: Papers highlighting key challenges. Studies that fall into more than one challenge are highlighted with square brackets.
@@
1 Error-proneness No. 8, 48,[25, 47, 26, 62, 27, 49, 59]
2 Imbalanced data No. [63, 85, 86, 50, 87, 68, 69, 70, 72, 89, 90, 91, 55, 93, 94, 96, 75, 101, 102, 105, 78, 65, 71, 88, 28, 97, 56, 57, 100], 84, 104, 106, 80, 60
3 Generalizability No. [26, 62, 27, 49, 25, 64, 85, 50, 68, 69, 72, 86, 89, 55, 93, 94, 96, 75, 101, 102, 105, 59, 108, 78], 67, 83, 110, 95, 73, 76, 31, 81, 109
@@
Error proneness
The first challenge is the dependence on manual analysis, which introduces the potential for human error and inconsistency 48. Since the process relies on human judgment, it is prone to inconsistencies in labeling, as different individuals may interpret and categorize SATD in different ways 47. This subjectivity in analysis can lead to varying outcomes in tasks such as pattern recognition and feature engineering, thereby reducing the reliability of the process.
The variability in human interpretation introduces a lack of uniformity in the labeled datasets, which can significantly affect the effectiveness of ML models that rely on this data for training. When models are trained on datasets that contain inaccuracies or inconsistencies, their predictive accuracy and generalization capabilities suffer. This challenge becomes even more pronounced when scaling up to larger datasets, where consistent and precise labeling is critical to maintaining the quality of the training data. Without robust and reliable analysis, models may struggle to handle the increased complexity of large datasets, resulting in reduced performance and misclassification of SATD instances.
@@
In most software projects, SATD comments represent only a small portion of the total code comments, resulting in a significant class imbalance. This imbalance can adversely affect the performance of SATD identification models, as they may focus on non-SATD comments that dominate the dataset. As a result, the models may struggle to accurately identify the relatively rare SATD instances, often prioritizing the majority class and overlooking the minority class 28.
This imbalance frequently leads to a high number of false negatives 117, in which the model fails to detect SATD instances, thereby undermining its effectiveness. Since ML models generally find it difficult to learn patterns from underrepresented classes 118, the scarcity of SATD examples makes it challenging for these models to develop a robust understanding of SATD characteristics. Consequently, the model’s ability to identify and classify SATD instances diminishes, particularly when faced with a large volume of non-SATD items. Addressing this imbalance is critical to improving the precision and reliability of SATD identification models, as it ensures that models are not biased toward detecting the majority class while neglecting crucial SATD instances.
@@
One of the primary challenges in identifying SATD is its generalizability. Generalizability refers to the extent to which SATD identification models maintain their accuracy and effectiveness when applied across different software projects and domains. This challenge arises from the diverse nature of software projects, which often exhibit significant variation in coding practices, domain-specific terminology, and architectural designs 54.
Models trained on data from one specific project tend to be tailored to the characteristics of that project, limiting their performance when applied to other projects with different practices. For example, the way SATD is expressed in comments may vary across projects due to differences in the development team’s language, domain-specific jargon, or coding style conventions 67, 49. These project-specific factors can result in models that are highly effective within their training environment but struggle to generalize to other contexts, especially when encountering unfamiliar coding practices or new domains.
@@
Another primary challenge in the SATD identification process, particularly in ML and DL-based models, is the lack of explainability. This issue arises because many advanced models, such as neural networks and DL architectures, function as “black boxes,” meaning their internal decision-making processes are not easily interpretable by users 119. These models generate predictions without providing transparent explanations about the underlying reasoning for classifying an instance as SATD, non-SATD, or even specific types of SATD.
@@
False negatives are another concern in SATD detection. They occur when SATD items are misclassified as non-SATD, resulting in undetected technical debt in the codebase. This occurs when models fail to identify instances in which developers have either implicitly or explicitly admitted technical debt 48. As a result, undetected technical debt can linger in the system, leading to long-term code degradation, higher maintenance costs, and an increased risk of introducing more bugs and system failures 13, 120.
@@
Another primary challenge in the context of machine learning and supervised model-based models is the issue of reproducibility 54, 74. These models often involve complex architectures with numerous parameters that must be fine-tuned for optimal performance. However, many studies fail to provide sufficient details on the hyperparameters, training configurations, or the specific processes used to achieve their results. Additionally, the lack of publicly available code exacerbates reproducibility issues, as other researchers cannot replicate the exact experimental setup. This makes it difficult to reproduce results across different studies and datasets, raising concerns about the reliability and validity of these techniques.
@@
Another challenge associated with SATD detection is the high computational costs of many advanced models, particularly DL and Transformer techniques. These models require significant resources during training, including high-performance GPUs, large memory capacities, and extended computation time. This high computational cost poses several obstacles for researchers and practitioners who aim to implement or improve SATD identification methods.
Additionally, the extended computation time required for training DL models presents a challenge in terms of practicality. SATD identification models often require retraining or fine-tuning to adapt to new projects or datasets, and the time-consuming nature of this process can hinder agile software development practices. In scenarios where rapid feedback is essential, such as continuous integration and deployment environments, the lengthy training times of DL models can become a bottleneck, delaying the detection of technical debt and the subsequent mitigation efforts.
@@
Furthermore, the overlap of features among different types of technical debt, such as code debt, design debt, and requirement debt, presents another challenge 28. This overlap makes distinguishing between different debt types difficult, leading to potential misclassification. The ambiguity in distinguishing between SATD and other debt types can reduce the accuracy of identification models. If features that are common to multiple debt types are not carefully handled, models may struggle to identify SATD correctly.
@@
Finally, tagging and comment relevance are a critical challenge in identifying SATD. Developers often use tags such as “TODO,” “FIXME,” or “XXX” in code comments or other textual artifacts to mark areas that require attention or indicate temporary fixes. While these task-related tags are commonly employed as indicators of SATD, they are not always relevant to technical debt. In many cases, tags such as “TODO” or “FIXME” can serve as simple reminders or non-technical debt items, thereby creating noise in the identification process. Moreover, not all SATD instances are explicitly marked with such tags, further complicating the identification process.
The results highlight significant trends in SATD detection, demonstrating an evolution from heuristic-based methods to advanced machine learning and deep learning techniques. However, these findings raise important questions regarding the applicability, generalizability, and practical adoption of these approaches. The following section presents a critical discussion of our results, compares them with the existing literature, and outlines their implications for future research and industry practice.
Answer to RQ5 We identified nine major challenges in detecting SATD, including manual analysis errors, data imbalance, and limited generalizability and explainability in advanced models. Other challenges include false negatives, reproducibility risks, computational costs, overlapping features among debt types, and the relevance of tagging and comments, all of which hinder effective SATD identification.